You have one type of candy that sells for $2.00/lb and another type of candy that sells for $8.00/lb. You would like to have 30 lbs of a candy mixture that sells for $3.90/lb. How much of each candy will you need to obtain the desired mixture?
You will need
? lbs of the cheaper candy
and
? lbs of the expensive candy.

Answer by jorel1380(3719)

m+n=30
2m+8n=30(3.90)=117
2m+2n=60
6n=57
n=57/6
m=123/6
You need 123/6 lbs. of $2 candy and 57/6 lbs. of $8 candy for your mix..